OpenStudy (anonymous):

How does a quartic function with 4, equal real roots look like?

hartnn (hartnn):

since they are equal roots, let them all be a. where a is real. now there are 4 roots, so (x-a)^4=0 it will look like this.

OpenStudy (anonymous):

Alright, how does the graph look like now?

OpenStudy (anonymous):

No longer a quadratic, but: \[f(x) = (x+1)^4\] Is one example. The general form is \[f(x) = (x+a)^4\] where "a" is any constant, and the solution will be -a.

OpenStudy (anonymous):

The graph will look like this: http://www.wolframalpha.com/input/?i=%28x%2B1%29%5E4 (For my example of \[f(x) = (x+1)^4\]) Note that for any function with multiple equivilant roots, the graph will "touch" the y-axis without crossing it.

hartnn (hartnn):

the graph of f(x)=(x-a)^4 will have one x-intercept of (a,0).

OpenStudy (anonymous):

* The graph will touch without crossing the x-axis for any even number of equivilant roots. Sorry for the confusion.

OpenStudy (anonymous):

Does the graph get wider as the exponent n increases for a function with n roots?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Ugh if i can only give medals to both of you instead of just one